OpenStudy (anonymous):

7x-3=28 -- --- x-4 x-4

OpenStudy (anonymous):

looks like you should try cross multiplication, are you sure thats an equal sign there?

OpenStudy (anonymous):

\[\left(\begin{matrix}7 \\ x-4\end{matrix}\right)- 3 =\left(\begin{matrix}28 \\ x-4\end{matrix}\right)\]

OpenStudy (anonymous):

okay, i got you

OpenStudy (anonymous):

then the first thing we need to do is get a common denomianto for the left side, can you do that?

OpenStudy (anonymous):

yeah i got